To treat an illness costed $10,000 (dollars in 2019) ten years ago. A new technology to treat this illness costs $10,000 (dollars in 2019) now. Do you think that there is zero inflation in the treatment specific for this illness?
Answer #1

There is an inflation in the price required to treat illness. The price required to treat illness 10 years ago, is less than $10000, but the price when adjusted for 2019, becomes $10000. Today, the cost of treatment is $10000 in 2019 dollar. Hence, when there is actual or current dollars, then there is a price difference and inflation exists.

For example,

Let,

CPI 10 years ago = 50

CPI in 2019 = 100

Then,

If price 10 years ago, is $5000, then as per the 2019 CPI adjustment, price = 5000(100/50) = $10000

Today's price at 2019 dollar = $10000

So, inflation rate in 10 years = (100-50)/50 = 100%

Above is the illustration that shows the presence of inflation as it is calculated on current or actual cost.